# Cell as Point: One-Stage Framework for Efficient Cell Tracking

### v1

@@ Abstract
Cellular activities are dynamic and intricate, playing a crucial role in advancing diagnostic and therapeutic techniques, yet they often require substantial resources for accurate tracking. Despite recent progress, the conventional multi-stage cell tracking approaches not only heavily rely on detection or segmentation results as a prerequisite for the tracking stage, demanding plenty of refined segmentation masks, but are also deteriorated by imbalanced and long sequence data, leading to under-learning in training and missing cells in inference procedures. To alleviate the above issues, this paper proposes the novel end-to-end CAP framework, which leverages the idea of regarding Cell as Point to achieve efficient and stable cell tracking in one stage. CAP abandons detection or segmentation stages and simplifies the process by exploiting the correlation among the trajectories of cell points to track cells jointly, thus reducing the label demand and complexity of the pipeline. With cell point trajectory and visibility to represent cell locations and lineage relationships, CAP leverages the key innovations of adaptive event-guided (AEG) sampling for addressing data imbalance in cell division events and the rolling-as-window (RAW) inference method to ensure continuous tracking of new cells in the long term. Eliminating the need for a prerequisite detection or segmentation stage, CAP demonstrates strong cell tracking performance while also being 10 to 55 times more efficient than existing methods. The code and models will be released.
@@ 1 Introduction
Indeed, several existing works and platforms [21] have demonstrated relatively effective performance in achieving automated cell tracking. The taxonomies of strategies typically involve segmentation followed by linking (SegLnk) [19, 45], segmentation and linking (Seg&Lnk) [14, 35, 40], or even detection prior to the segmentation and linking stages (DetSegLnk [49] or DetLnkSeg [36, 20][43]. Also, frameworks [24, 25] often incorporate isolated data-preprocessing stages to represent cell tracking data. These multi-stage cell tracking pipelines can limit the overall performance of the framework, constrained by the errors that occur at earlier stages and are accumulated and amplified at subsequent stages [37]. Additionally, a multi-stage framework involving both detection/segmentation and tracking not only requires the tracking ground-truth but also demands additional labels, i.e., segmentation masks, to train the model. It is highly labor-intensive and time-consuming to produce the data annotation, rendering such methods impractical for clinical applications [72].
@@
To overcome the data challenge for existing multi-stage frameworks, we propose a novel one-stage cell tracking framework, Cell as Point (CAP), as demonstrated in Fig. 1 (a), that treats each cell as a point to jointly track numerous cell points within a sequence. The one-stage CAP framework not only benefits from the efficient prediction procedure through a straightforward pipeline, shown in Fig. 1 (b), but also from its low requirement for data annotation, solely relying on tracking ground-truth (coarse masks and cell acyclic graphs) to represent the position and status of each cell. Specifically, by proposing cell point trajectory and visibility to represent the cell tracking data within one-stage framework, CAP offers an elegant solution to alleviate the issues of the complicated data-preprocessing and the dilemma of representation of cell decay, which existing methods struggle to cope with. Another highlight of CAP is the creation of adaptive event-guided (AEG) sampling method and rolling-as-window (RAW) inference approach. Specifically, AEG sampling is a temporal data sampling strategy for cell microscopy imaging sequence data to remedy the imbalance of cell division occurrences in the dataset; RAW is an inference approach to capture and track new cells within a long sequence in an efficient way. To summarize, our key contributions include:
@@
  • We propose a novel CAP cell tracking framework that is highly efficient in terms of inference time and data annotation requirements, enabling one-stage tracking that bypasses detection/segmentation to observe and investigate cellular behaviors.
@@
  • We address prevalent challenges in the cell tracking field by proposing AEG sampling to resolve imbalanced cell division occurrence, representing complex cell events for training purposes through cell point trajectory and visibility concepts within a one-stage framework, and employing a RAW inference approach to capture and track new cells in long sequences.
@@
  • Our one-stage approach, CAP, achieves comparable or higher performance compared to multi-stage methods while alleviating the need for detection or segmentation.
@@ 2.1 Cell Tracking
Cell tracking is a crucial task in biological imaging that typically involves the identification and temporal association of cell instances across image frames. Traditionally, this process includes two main steps: cell segmentation (detection) and cell linking [13]. The segmentation step identifies individual cells in each frame, while the linking step establishes temporal associations between these cells across frames, sometimes capturing complex dynamics such as cell division and migration [40, 43]. Nowadays, instead of traditional approaches such as Nearest Neighbors linking [56], Bayesian filters [54, 12, 26, 67, 27, 61], Viterbi [42, 41], and graph-based matching [3, 60, 2, 39], the advent of deep learning has transformed the landscape of cell tracking, leveraging advanced neural architectures for segmentation and tracking tasks. Deep learning approaches often utilize convolutional neural networks (CNNs) and recurrent neural networks (RNNs) to perform these tasks [55], employing complex models that combine multiple stages like detection [10, 8, 9, 38, 74, 23, 46, 65, 24, 25], segmentation [68, 47, 7, 1, 40, 57, 5, 6, 52], and tracking into the frameworks. Several contemporary methods [50, 48, 23] aim to integrate contextual cell data into tracking models. ConvGRU [48] is introduced by Payer et al., capturing local cellular and inter-frame features. However, this method requires extensive annotated training data for all cell regions, making it labor-intensive. Motion and Position Map (MPM) is capable of representing detection and association at once, which is proposed by Hayashida et al. [24], marking an advancement in simplifying the pipeline for cell tracking. However, it still requires multiple stages, first generating the MPM followed by subsequent detection and tracking procedures.
@@ 2.2 Point Tracking
In the field of natural images, recent end-to-end video point-tracking frameworks have achieved commendable performance. Notably, TAP-Vid [15] pioneered tracking any point across video frames, establishing a benchmark, and claiming a baseline method. However, this method does not support tracking occluded points, limiting its applicability in cell-tracking scenarios. Even though PIPs [22] and PointOdyssey [73] methods take measures to address the issue of point occlusion in long sequences by tracking selected points using a fixed sliding window and resuming tracking from the last visible frame, these methods struggle with point tracking independently, which may not capture the dependencies between multiple points. OmniMotion [64] accounts for estimated correspondences within a canonical space by optimizing a volumetric representation for video sequences. Although innovative, this method involves test-time optimization, making it computationally expensive, which is less viable for clinical applications. CoTracker [32] represents significant advancements in the field of point tracking. It addresses the inaccuracies led by independent tracking among points by incorporating incremental point coordinates updates based on 4D correlation volumes, first proposed by RAFT [59] constructed for all pairs of points from point/object features. CoTracker3 [31] is advanced by involving a semi-supervised training methodology to reduce reliance on synthetic datasets and utilizing pseudo-labeling techniques to generate training labels from unannotated real-world videos.
@@
Although the performance of cell tracking may benefit from investigating the interdependencies between each cell, inspired by joint tracking [32], the unique properties of cell behaviors, i.e., cell mitosis and apoptosis, have not been considered in the model construction. Specifically, cells are not only relative with position and occlusion but also potentially associated with their ancestors due to the regulation of cellular life decay and division, triggering another issue: missing tracking for new daughter cells. Moreover, regarding the publicly available cell tracking data [45, 43, 62, 33], the diversity of cell types results in significant variations in cell division rates and sequence length across different datasets categorized by specimens. Imbalanced data [34] can lead to the omission of complete division events (from the death of the mother cell to the birth of the daughter cell) during random sampling with a low division rate; long sequences typically require more time or computational resources to predict cell locations and lineages at once. Thus, existing methods potentially result in the under-learning of division cases in the model training process and a higher resource requirement in the inference procedure.
Therefore, the uniqueness of this time-lapse cell sequential image data poses the most significant challenge in developing a one-stage cell tracking framework to effectively track cells jointly.
@@ 3 Cell as Point (CAP)
Figure 2: Cell Point Trajectory and Visibility. M, D1, and D2 denote mother cell, daughter cell 1, daughter cell 2. Three valid trajectory and visibility combinations: (a) M has not divided; (b) M has divided, and only D1 occurs in the $t$-th frame; (c) M has divided, and both D1 and D2 occur in the $t$-th frame.
Cell tracking aims to identify the characteristics of cell migration and division, thereby fully automating the prediction of specific cell position and division status changes in each frame of a time-lapse microscopy video. In our framework, CAP, every single cell is a point, tracking all cell points throughout the duration of a frame sequence $S=(I_{t})_{t=1}^{T}$ containing $T$ grayscale cell microscopy images $I_{t}\in\mathbb{R}^{H\times W}$. CAP primarily utilizes and learns the dependencies between cell point trajectories and visibilities to predict cell locations and lineages. The cell point trajectory is constituted by the $N$ cell centroid point locations $L^{i}_{t}=(x^{i}_{t},y^{i}_{t})\in\mathbb{R}^{2}$, $i=1,\dots,N$, $t=1,\dots,T$, where $t$ is the time of a frame sequence. The cell point visibility of the cell $V^{i}_{t}\in\{0,1\}$ indicates whether the $i$-th cell point appears at the location $L^{i}_{t}$ in the frame at time $t$; in other words, whether the cell has divided into new cells and disappeared.  Fig. 2 demonstrates the combination of cell point trajectory and visibility, representing various states of cells, such as migration, mitosis, apoptosis, and cell lineages (with two potential daughter cells). This simple yet effective representation method will efficiently aid the CAP framework in achieving one-stage efficient cell tracking.
@@ 3.1 Adaptive Event-Guided (AEG) Sampling
To facilitate the effective training of models to understand the biological behavior of cell division while mitigating data imbalance [70], we propose the Adaptive Event-Guided (AEG) sampling technique to balance the occurrences of cell division in a sequence of frames the model sees. In order to reduce operational costs and compromise overfitting [4], instead of seeing the entire training sequence in a single instance, the model learns a continuous sequence of $T_{s}$ frames, where $0<T_{s}<T$, with the initial frame index determined using AEG for each iteration. The AEG method leverages the intrinsic dynamics of cell division within the frame sequences to guide the sampling process, enhancing the representational diversity of the training dataset and ensuring the model’s robustness to variations in cell behavior. Technically, as data loading iterates through the whole dataset by detecting the complete cell division processes (from the death of parent cell to the birth of daughter cell), an anchor is added at each time cell division occurs $\mathbb{T}_{anchors}$. When the probability $P_{AEG}$ indicates the application of the AEG method, our CAP framework enables AEG sampling applied to selecting $T_{s}$ as illustrated in  Fig. 3; Otherwise, the starting frame $t_{starting}$ of $T_{s}$ is randomly located in the range of $(0,T-T_{s})$.
@@
Figure 3: AEG Sampling. We propose integrating AEG sampling on the sequence of frames to enable the model to learn a complete and effective cell division process.
@@ 3.2 Cell Joint Tracking
By regarding each cell as a point and leveraging the two-dimensional feature across time and a set of tracked cell points, we implement an effective and efficient method for cell tracking via the correlation between cell point trajectories and visibilities. Specifically, the training process of CAP as a transformer $\Theta$ neural network [63], where $\Theta:G\mapsto O$. In order to improve the estimated cell trajectories and visibilities along with $M$ iterations, the trajectories and visibilities of $i$ points (cells) in a $t$-frame sequence are encoded into a grid of input tokens $G^{i}_{t}$. $O^{i}_{t}$ is the corresponding output of $\Theta$ for representing the refined trajectories and visibilities. The following sections elaborate on the transformer formulation visualized in Fig. 4.
Figure 4: Cell Joint Tracking Structure. For a sequence $T_{s}$, evolving from initialized to estimated cell point trajectories and visibilities requires $M$ updates on cell point trajectory $\hat{L}$ and tracking feature $F$ with cell point visibility $\hat{V}$ computed at the last ($M$-th) time update.
@@
To initialize the tracking feature $F$ for the part of input tokens $G$, we first generate image features (feature map) for each frame in the training sequence, from which the tracking features are distilled according to cell point locations. Specifically, the feature extractor is a convolutional neural network (CNN) trained end-to-end in the framework for computational efficiency. Leveraging the CNN feature extractor, the feature maps with $dim$ dimensions are generated from $I_{t}$, $\phi(I_{t}):\mathbb{R}^{3\times H\times W}\mapsto\mathbb{R}^{dim\times\left(\left\lfloor\frac{H}{k}\right\rfloor\times\left\lfloor\frac{W}{k}\right\rfloor\right)}$, where $k=4$ for computational efficiency, and the features are downscaled by average pooling with strides $s$ for multi-scaled features.
The tracking features $F^{i}_{t=1}$ for $i$ are captured by bilinear sampling from the location of cell points on the initial frame $L_{t=1}$ on the feature map $\phi(I_{t=1})$. By broadcasting the tracking features in $t=1$ frame for all training frames $T_{s}$ into shape $(T_{s},N,dim)$, the tracking feature vector $F^{i}_{t}\in\mathbb{R}^{dim}$ is initialized for the transformer input, which is refined and updated with the training process.
@@
To achieve the objective of associating cell point trajectories with a sequence of frames, CAP employs correlation volume $C^{i}_{t}\in\mathbb{R}^{S}$, where $S=4$, proposed in RAFT [59]. These correlation volumes $C^{i}_{t}$ are derived by assessing the relationship between the tracking features $F^{i}_{t}$ and the image features $\phi(I_{t})$ in proximity to the estimated cell point location $\hat{L}^{i}_{t}$ of the trajectory on the $t$-th frame. Technically, each $C^{i}_{t}$ is the average over the correlation response in a $2^{s}\times 2^{s}$ grid:
@@
where $s=1,\dots,S$ are the feature scales and $\delta\in\mathbb{Z}^{2}$, $\|\delta\|_{\infty}\leq\Delta$ are offsets around cell points, $\hat{x}$ and $\hat{y}$ are the estimated location of the cell point $\hat{L}_{t}^{i}$. This means that the value at $[C_{t}^{i}]_{\delta}^{s}$ can be computed as the inner product between the feature vector $F^{i}_{t}$ and $\phi_{s}(I_{t})$ at $\hat{L}_{t}^{i}$ pooled with kernel size $2^{s}\times 2^{s}$. All different scaled features are sampled by using bilinear interpolation at non-integer locations and border padding from the image feature map $\phi_{s}(I_{t})$.
Input and Output Tokens.
The input tokens express the concatenation of features: location $\hat{L}^{i}_{t}-\hat{L}^{i}_{1}$, visibility $\hat{V}^{i}_{t}$, tracking features $F^{i}_{t}$, and correlation volume of cell trajectories $C^{i}_{t}$. As for the output tokens $O(\Delta\hat{L},\Delta F)$, contain updates for location and refined tracking features:
$O^{i}_{t}=(\Delta\hat{L}^{3i}_{t},\Delta F^{i}_{t}),$ (2)
where CAP estimates $3$ locations for each cell, following the structure in Fig. 2: the cell’s own location and the locations of two potential daughter cells. This is because any cell division that happened in the event would have been that of the parent cell into two daughter cells [29]. Thus, CAP represents $L$ and $V$ through a $3$-column format to reserve space for each cell’s progeny.
@@
Utilizing the approach of applying the transformer $\Theta$ iteratively for a sequence $T_{s}$, the update operator in CAP produces and progressively improves the cell locations estimation $\hat{L}^{3i}_{t}$. Considering $M$ times estimating the cell locations, the update of the output computes $O(\Delta\hat{L},\Delta F)$, which is applied to the next iteration:
@@
where $m=1,\dots,M$ and $m=1$ denotes initialization. At the last iteration of the update $M$, cell point visibility $V^{3i}_{t}$ for locations estimated $\hat{L}^{3i}_{t}$ is computed by $\hat{V}^{(M)}=\sigma(WF^{(M)})$, where $\sigma$ represents the sigmoid activation function and $W$ is a learned matrix of weights. By investigating the $3$ visibilities $V^{3i}_{t}$, CAP tells if the cell division occurred on cell $i$ in the $t$-th frame.
Supervision.
We supervised our transformer $\Theta$ on the trajectory and visibility between predicted and ground truth for every semi-overlapping window. The trajectory loss function $\mathcal{L}_{tra}$ is location regression summed over the update iterations, and the visibility loss function $\mathcal{L}_{vis}$ is the cross entropy of binary visibility options:
@@
where $M$ is the number of times iteratively estimating the cell locations with updated $\mathcal{L}_{tra}$ for a sequence, and $\mathcal{L}_{vis}$ is updated at the $M$-th iteration; $J$ is the total number of $T_{s}$ whose trajectory starts in the middle of the previous $T_{s}$.
3.3 Rolling-as-Window (RAW) Inference
In the field of cell tracking, the inference part plays a crucial role in generating the final tracking mask and in documenting cell lineage relationships based on predicted cell locations and division events. In existing works [57, 52, 65, 40, 6], the inference and post-processing stages are mostly explicitly tailored to different datasets, which vary in sequence lengths, which may hinder broad application to other datasets. In contrast, our proposed inference approach, rolling-as-window (RAW) in the CAP, not only simplifies and unifies these processes but also proves effective across varying lengths of frame sequences, facilitating high-quality inference regardless of computational resources. Within our model, the mechanism of iterative updates assists in maintaining the quality of inference results. It ensures that the output remains consistent and high-quality while utilizing a rolling loop that crops the yet-to-be-inferred clip from the whole sequence into a window for inference.

### v2

@@ Abstract
Conventional multi-stage cell tracking approaches rely heavily on detection or segmentation in each frame as a prerequisite, requiring substantial resources for high-quality segmentation masks and increasing the overall prediction time. To address these limitations, we propose CAP, a novel end-to-end one-stage framework that reimagines cell tracking by treating Cell as Point. Unlike traditional methods, CAP eliminates the need for explicit detection or segmentation, instead jointly tracking cells for sequences in one stage by leveraging the inherent correlations among their trajectories. This simplification reduces both labeling requirements and pipeline complexity. However, directly processing the entire sequence in one stage poses challenges related to data imbalance in capturing cell division events and long sequence inference. To solve these challenges, CAP introduces two key innovations: (1) adaptive event-guided (AEG) sampling, which prioritizes cell division events to mitigate the occurrence imbalance of cell events, and (2) the rolling-as-window (RAW) inference strategy, which ensures continuous and stable tracking of newly emerging cells over extended sequences. By removing the dependency on segmentation-based preprocessing while addressing the challenges of imbalanced occurrence of cell events and long-sequence tracking, CAP demonstrates promising cell tracking performance and is 10 to 55 times more efficient than existing methods. The code and model checkpoints will be available soon.
@@ 1 Introduction
Indeed, several existing works and platforms [28, 25, 26] have demonstrated relatively effective performance in achieving automated cell tracking. The taxonomies of strategies typically involve segmentation followed by linking (SegLnk) [24, 54], segmentation and linking (Seg&Lnk) [19, 43, 49], or even detection prior to the segmentation and linking stages (DetSegLnk [61] or DetLnkSeg [44, 27][52]. Also, frameworks [31, 32] often incorporate isolated data-preprocessing stages to represent cell tracking data. These multi-stage cell tracking pipelines can limit the overall performance of the framework, constrained by the errors that occur at earlier stages and are accumulated and amplified at subsequent stages [45]. Additionally, a multi-stage framework involving both detection/segmentation and tracking not only requires the tracking ground-truth but also demands additional labels, i.e., high-quality segmentation masks, to train the model. It is highly labor-intensive and time-consuming to produce the data annotation, rendering such methods impractical for clinical applications [86].
@@
To overcome the data challenge for existing multi-stage frameworks, we propose a novel one-stage cell tracking framework, Cell as Point (CAP), as demonstrated in Fig. 1 (a), that treats each cell as a point to jointly track numerous cell points within a sequence. The one-stage CAP framework not only benefits from the efficient prediction procedure through a straightforward pipeline, shown in Fig. 1 (b), but also from its low requirement for data annotation, solely relying on tracking ground-truth (coarse masks and cell acyclic graphs) to represent the position and status of each cell. Specifically, by proposing cell point trajectory and visibility to represent the cell tracking data within a one-stage framework comprehensively, CAP offers an elegant solution to alleviate the issues of the complicated data-preprocessing and the dilemma of representation of cell decay. Another key innovation of CAP is the introduction of adaptive event-guided (AEG) sampling method and rolling-as-window (RAW) inference approach. Specifically, AEG sampling is a temporal data sampling strategy for cell microscopy imaging sequence data to remedy the imbalance of cell division occurrences in the dataset; RAW is an inference approach to capture and track new cells within a long sequence in an efficient way. To summarize, our key contributions include:
@@
  • We propose a novel CAP cell tracking framework that is highly efficient in terms of inference time and data annotation requirements, enabling one-stage tracking that bypasses detection/segmentation to investigate cellular behaviors.
@@
  • We address prevalent challenges in the cell sequence data by proposing AEG sampling to resolve imbalanced cell division occurrence, representing complex cell events for training purposes through cell point trajectory and visibility concepts within a one-stage framework, and employing a RAW inference approach to capture and track new cells in long sequences.
@@
  • Our one-stage approach, CAP, achieves promising performance compared to existing multi-stage methods while alleviating the need for detection or segmentation.
@@ 2.1 Cell Tracking
Cell tracking is a crucial task in biological imaging that typically involves the identification and temporal association of cell instances across image frames. Traditionally, this process includes two main steps: cell segmentation (detection) [1, 79, 55, 77] and cell linking [18]. The segmentation step identifies individual cells in each frame, while the linking step establishes temporal associations between these cells across frames, sometimes capturing complex dynamics such as cell division and migration [49, 52]. Nowadays, instead of traditional approaches such as Nearest Neighbors linking [67], Bayesian filters [66, 16, 33, 81, 34, 73], Viterbi [51, 50], and graph-based matching [5, 72, 3, 48], the advent of deep learning has transformed the landscape of cell tracking, leveraging advanced neural architectures for segmentation and tracking tasks. Deep learning approaches often employ complex models, e.g., CNN, RNN, GNN, and transformer, to combine multiple stages like detection [12, 10, 11, 46, 88, 30, 56, 78, 31, 32, 38], segmentation [82, 59, 9, 2, 49, 68, 7, 8, 64, 58, 17, 13, 57, 14, 71], and tracking into the frameworks. Several contemporary methods [62, 60, 30] aim to integrate contextual cell data into tracking models. ConvGRU [60] is introduced by Payer et al., capturing local cellular and inter-frame features. However, this method requires extensive annotated training data for all cell regions, making it labor-intensive. Motion and Position Map (MPM) is capable of representing detection and association at once, which is proposed by Hayashida et al. [31], marking an advancement in simplifying the pipeline for cell tracking. However, it still requires multiple stages, first generating the MPM followed by subsequent detection and tracking procedures.
@@ 2.2 Point Tracking
In the field of natural images, recent end-to-end video point-tracking frameworks have achieved commendable performance. Notably, TAP-Vid [20] pioneered tracking any point across video frames, establishing a benchmark, and claiming a baseline method. However, this method does not support tracking occluded points, limiting its applicability in cell-tracking scenarios. Even though PIPs [29] and PointOdyssey [87] methods take measures to address the issue of point occlusion in long sequences by tracking selected points using a fixed sliding window and resuming tracking from the last visible frame, these methods struggle with point tracking independently, which may not capture the dependencies between multiple points. OmniMotion [76] accounts for estimated correspondences within a canonical space by optimizing a volumetric representation for video sequences. Although innovative, this method involves test-time optimization, making it computationally expensive and less viable for clinical applications. CoTracker [40] represents significant advancements in the field of point tracking. It addresses the inaccuracies led by independent tracking among points by incorporating incremental point coordinates updates based on 4D correlation volumes, first proposed by RAFT [70] constructed for all pairs of points from point/object features. CoTracker3 [39] is advanced by involving a semi-supervised training methodology to reduce reliance on synthetic datasets and utilizing pseudo-labeling techniques to generate training labels from unannotated real-world videos.
@@
However, the trackers designed for general-purpose applications are ineffective for cell tracking because the unique properties of cell behaviors, i.e., cell mitosis and apoptosis, have not been considered in the model construction. Specifically, cells are not only relative with position and occlusion but also potentially associated with their ancestors due to the regulation of cellular life decay and division. Therefore, they cannot simultaneously represent or predict both cell positions and division states with their ancestors or children. Moreover, regarding the publicly available cell tracking data [54, 52, 74, 41, 4], the diversity of cell types results in significant variations in cell division rates and sequence length across different datasets categorized by specimens. Imbalanced cell event occurrence [42] in a sequence can lead to the omission of complete division events (from the death of the mother cell to the birth of the daughter cell) during random sampling with a low division rate; the inference for long sequences typically requires more time or computational resources to predict cell locations and lineages at once. Thus, existing methods potentially result in the under-learning of division cases in the model training process and a higher resource requirement in the long-sequence inference procedure. Thus, the uniqueness of this time-lapse cell sequential image data poses the most significant challenge in developing a one-stage cell tracking framework to effectively track cells jointly.
@@ 3 Cell as Point (CAP)
Figure 2: Cell Point Trajectory and Visibility. M, D1, and D2 denote mother cell, daughter cell #1, and daughter cell #2. (x,y) in $L_{t}$ represent the location coordinates, and 00 or $1$ in $V_{t}$ respectively represent non-existing and existing cells. Three valid trajectory and visibility combinations: (a) M has not divided; (b) M has divided, and only D1 occurs in the $t$-th frame; (c) M has divided, and both D1 and D2 occur in the $t$-th frame.
Cell tracking aims to identify the characteristics of cell migration and division, thereby fully automating the prediction of specific cell position and division status changes in each frame of a time-lapse microscopy video. In our framework, CAP, every single cell is a point, tracking all cell points throughout the duration of a frame sequence $S=(I_{t})_{t=1}^{T}$ containing $T$ grayscale cell microscopy images $I_{t}\in\mathbb{R}^{H\times W}$. Cell point trajectories and visibilities are proposed to be utilized in the CAP framework to represent and predict cell locations and lineages. The cell point trajectory is constituted by the $N$ cell centroid point locations $L^{i}_{t}=(x^{i}_{t},y^{i}_{t})\in\mathbb{R}^{2}$, $i=1,\dots,N$, $t=1,\dots,T$, where $t$ is the time of a frame sequence. The cell point visibility of the cell $V^{i}_{t}\in\{0,1\}$ indicates whether the $i$-th cell point still exists in the frame sequence at time $t$; in other words, if the cell is non-existing or born due to mitosis and apoptosis. Since any cell division that happened in the event would have been that of the parent cell into two daughter cells [36]Fig. 2 demonstrates the combinations of cell point trajectory and visibility, representing various states of cells, i.e., migration, mitosis, apoptosis, and cell lineages (with two potential daughter cells). This simple yet effective representation method will efficiently aid the CAP framework in achieving one-stage efficient cell tracking.
@@ 3.1 Adaptive Event-Guided (AEG) Sampling
Figure 3: Overview of CAP framework. The training sequence $T_{s}$ is sampled from the entire frame sequence using AEG strategy. For each $T_{s}$, the cell point trajectories and visibilities are iteratively refined, starting from their initialization. For iteration $1,\dots,M$, updates are made to cell point trajectory $\hat{L}$ and tracking feature $F$ with cell point visibility $\hat{V}$ computed at the final ($M$-th) update. This figure illustrates a single iteration of the optimization process.
To reduce operational costs and compromise overfitting [6], instead of seeing the entire training sequence in a single instance, the model learns a continuous sub-sequence of $T_{s}$ frames for each time, where $0<T_{s}<T$. However, the imbalanced occurrence of cell division in the entire sequence would cause the model to detect cell division occurring rarely. To facilitate the effective training of models to understand the biological behavior of the completed cell division event while mitigating data imbalance [84, 47], we propose the Adaptive Event-Guided (AEG) sampling technique to balance the occurrences of cell division in a sequence of frames the model sees by determining the initial frame of $T_{s}$ learning a complete event.
The AEG method leverages the intrinsic dynamics of cell division within the frame sequences to guide the sampling process, enhancing the representational diversity of the training dataset and ensuring the model’s robustness to variations in cell behavior. Technically, as data loading iterates through the whole dataset by detecting the complete cell division processes (from the death of the parent cell to the birth of the daughter cell), an anchor is added at each time cell division occurs $\mathbb{T}_{anchors}$. The probability of applying AEG $P_{\text{AEG}}$ is adaptively calculated by $P_{\text{AEG}}=\frac{N_{\text{div}}\times T_{\text{div}}}{T}$, where $N_{\text{div}}$ denotes the total occurrence of cell division in the entire $T$-frame sequence, and $T_{\text{div}}$ denotes the duration of the completed cell division. When the random decimal $p<P_{\text{AEG}}$ indicates the usage of AEG, our CAP framework incorporates AEG sampling for selecting $T_{s}$, as illustrated at the top of Fig. 3; Otherwise, the starting frame of $T_{s}$ is randomly located in the $(0,T-T_{s})$ range.
@@ 3.2 Cell Joint Tracking
By regarding each cell as a point and leveraging the two-dimensional feature across time and a set of tracked cell points, we implement an effective method for cell tracking via the correlation between cell point trajectories and visibilities in one stage. The training process of CAP as a transformer $\Theta$ neural network [75], where $\Theta:G\mapsto O$. In order to improve the estimated cell trajectories and visibilities along with $M$ iterations, the trajectories and visibilities of $i$ points (cells) in a $t$-frame sequence are encoded into a grid of input tokens $G^{i}_{t}$. $O^{i}_{t}$ is the corresponding output of $\Theta$ for representing the refined trajectories and visibilities. The following sections elaborate on the transformer formulation visualized in Fig. 3.
@@
To initialize the tracking feature $F$ for the part of input tokens $G$, we first generate image features (feature map), $\phi(I_{t})$, for each frame in the training sequence, from which the tracking features, $F$, are distilled according to cell point locations. Specifically, the feature extractor is a convolutional neural network (CNN) trained end-to-end in the framework for computational efficiency. Leveraging the CNN feature extractor, the feature maps with $dim$ dimensions are generated from $I_{t}$, $\phi(I_{t})\in\mathbb{R}^{dim\times\left(\left\lfloor\frac{H}{k}\right\rfloor\times\left\lfloor\frac{W}{k}\right\rfloor\right)}$, where $k=4$ for computational efficiency, and the features are downscaled by average pooling with strides $s$ for multi-scaled features. The tracking features $F^{i}_{t=1}$ for $i$ are captured by bilinear interpolation from the location of cell points on the initial frame $L_{t=1}$ on the feature map $\phi(I_{t=1})$. The tracking feature vector $F^{i}_{t}\in\mathbb{R}^{dim}$ is initialized for the transformer input by broadcasting the tracking features in $t=1$ frame for all $T_{s}$ frames, which is refined and updated with $M$ iterations.
@@
To achieve the objective of associating cell point trajectories jointly, CAP employs correlation volume $C^{i}_{t}\in\mathbb{R}^{S}$, where $S=4$, proposed in RAFT [70]. These correlation volumes $C^{i}_{t}$ are derived by assessing the relationship between the tracking features $F^{i}_{t}$ and the image features $\phi(I_{t})$ in proximity to the estimated cell point location $\hat{L}^{i}_{t}$ of the trajectory on the $t$-th frame. Technically, each $C^{i}_{t}$ is the average over the correlation response in a $2^{s}\times 2^{s}$ grid:
@@
where $s=1,\dots,S$ are the feature scales, and $\delta$ is the offset around cell points, $(\hat{x},\hat{y})$ is the estimated location coordinate of the cell point $\hat{L}_{t}^{i}$. The value of $[C_{t}^{i}]_{\delta}^{s}$ is computed as the inner product between the feature vector $F^{i}_{t}$ and $\phi_{s}(I_{t})$ at $\hat{L}_{t}^{i}$ pooled with kernel size $2^{s}\times 2^{s}$.
@@
In the transformer-based interactive updates, the input tokens express the concatenation of features: location $\hat{L}^{i}_{t}-\hat{L}^{i}_{1}$, visibility $\hat{V}^{i}_{t}$, tracking features $F^{i}_{t}$, and correlation volume of cell trajectories $C^{i}_{t}$. As for the output tokens $O(\Delta\hat{L},\Delta F)$, contain updates for location and refined tracking features: $O^{i}_{t}=(\Delta\hat{L}^{3i}_{t},\Delta F^{i}_{t})$, CAP estimates $3$ locations for each cell, following the representation illustrated in Fig. 2: the cell’s location and the locations of two potential daughter cells. In this way, CAP reserves space for each cell’s progeny by representing $L$ and $V$ through a $3$-column format. $M$-times updates the output tokens by applying the transformer $\Theta$ iteratively for a sequence $T_{s}$ as shown in Fig. 3. Specifically, CAP utilizes transformer $\Theta$ with cross-trajectory/time attention to update both the cell point trajectory, $\hat{L}$, and the tracking feature, $F$:
@@
where $m=1,\dots,M$, and $m=1$ denotes initialization. At the last iteration of the update $M$, cell point visibility $V^{3i}_{t}$ for locations estimated $\hat{L}^{3i}_{t}$ is computed by $\hat{V}^{(M)}=\sigma(WF^{(M)})$, where $\sigma$ represents the sigmoid activation function and $W$ is a learned matrix of weights. By investigating the $3$ visibilities $V^{3i}_{t}$, CAP tells if the cell division occurred on the cell $i$ in the $t$-th frame.
Loss Function.
We supervised the transformer $\Theta$ on the trajectory and visibility between predicted and ground truth for every semi-overlapping window. The trajectory loss function $\mathcal{L}_{tra}$ is location regression summed over the iteration updates, and the visibility loss function $\mathcal{L}_{vis}$ is the cross entropy of binary visibility options:
@@
where $\mathcal{L}_{vis}$ is updated at the $M$-th iteration, and $J$ is the total number of $T_{s}$ whose trajectory starts in the middle of the previous $T_{s}$.